OpenStudy (anonymous):

why is sqrt7/7 = 1/sqrt7 ? Been reading through all the rules of exponents and radicals...and still can't piece together why this is....

OpenStudy (anonymous):

It's better to flip it backwards. 1/sqrt7 = sqrt7/7

OpenStudy (anonymous):

most people (teachers) don't want sqrt in the bottom (denominator)

OpenStudy (anonymous):

so you multiply 1/sqrt7 by sqrt7/sqrt7. The two sqrt7's turn into the 7. then the 1*sqrt7 turns into sqrt7 on the top.

OpenStudy (anonymous):

sorry i would write it out, but I'm not at my normal computer.

OpenStudy (anonymous):

oh wait I see now that you get this by rationalizing the denominator , this question i'm on though says to express 1/sqrt 7 as an exponent..is this possible?

hartnn (hartnn):

a can be written as \(\sqrt a.\sqrt a\) so 7 can be written as \(\sqrt 7.\sqrt 7\) hence \(\huge\frac{\sqrt 7}{7}=\frac{\sqrt 7}{\sqrt 7.\sqrt 7}=\frac{1}{\sqrt 7}\)

OpenStudy (anonymous):

yes, sqrt7 is the same thing as saying (7)^(1/2)

OpenStudy (anonymous):

ah I see now, so its basically what the book was talking about with the properties of nth roots or no?

OpenStudy (anonymous):

mmm yes I guess you could say that. 3rd root of 7 would be (7)^1/3

OpenStudy (anonymous):

sqrt's are basically fractional exponents.

OpenStudy (anonymous):

ah okay I get it now hah, Thank you!!

OpenStudy (anonymous):

np gl